OpenStudy (anonymous):

Which expression is the simplified form of sqrt k^17

OpenStudy (diyadiya):

\[\Large k^{17} = k^{16} \times k^1= k^{16} \times k\] \[\Large \sqrt{k^{16} \times k} = k^8 \sqrt{k}\]

OpenStudy (anonymous):

Thanks.

OpenStudy (diyadiya):

You're welcome!